1. A spinner is divided into six equally sized regions. Two of the regions are labeled 6, two of the regions are labeled 2, one region is labeled 1, and the last region is labeled 3. The spinner is spun 750 times.
(a) How many times is the spinner expected to land on 6?
(b) How many times is the spinner expected to land on a prime number?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52777)

.
1. A spinner is divided into six equally sized regions.
Two of the regions are labeled 6, two of the regions are labeled 2, one region is labeled 1,
and the last region is labeled 3. The spinner is spun 750 times.
(a) How many times is the spinner expected to land on 6?
(b) How many times is the spinner expected to land on a prime number?
~~~~~~~~~~~~~~~~


(a)  2/6, or 1/3 of 750, which is 250 times.    ANSWER



(b)  (2/6) related to regions labeled 2  PLUS  (1/6)  related to region labeled 3. 

           The total  is  2/6 + 1/6 = 3/6 = 1/2.


      1/2 of 750 is 375 times.    ANSWER

Solved.     Easy calculations.


------------


The error is just fixed.



Answer by greenestamps(13198)



For tutor @ikleyn...
Check your numbers for part (b)
(there is only 1 region labeled 3)